how much force would be required to produce 88 j of work when pushing a box 1.1meters at an angle of 10 degrees?
1
Expert's answer
2020-05-11T20:14:25-0400

The work is


"W=Fd\text{cos}\theta,"

where F, d, and θ are respectively force, distance, angle between the vector of force and the distance.

Express and find the force:


"F=\frac{W}{d\text{cos}\theta}=\frac{88}{1.1\cdot\text{cos}10^\circ}=81.2\text{ N}."